A Kenworth road tractor purchased by Rio Corporation cost $88,500. This item will be used for business 100% of the time. Accountants elected to take no section 179 deduction and utilize the special depreciation allowance of 30%. Prepare a depreciation schedule using the MACRS method and enter the accumulated depreciation at the end of year 3 as your answer. Round all dollar amounts to the nearest cent

(NEED ANSWER BY 11:30PM PLEASE)
Answer #1

Formula sheet

A B C D E F G H I J K
2
3 All vehicles fall under MACRS 5 year depreciation.
4 Cost basis 88500
5 Since special depreciation allowance for 30% has been taken, therefore new cost basis can be calculated as follows:
6 New Cost Basis =$88500 - 88500*30%
7 =$88500 - $26,550
8 =D4-D4*30%
9
10 Year Cost Basis x Depreciation rate Computed Depreciation Accumulated Depreciation
11 Year 1 88500 x 0.3 =D11*F11
12 =D11-G11 x 0.2 =D12*F12 =G12+G11
13 Year 2 =D12 x 0.32 =D13*F13 =G13+H12
14 Year 3 =D13 x 0.192 =D14*F14 =G14+H13
15 Year 4 =D14 x 0.1152 =D15*F15 =G15+H14
16 Year 5 =D15 x 0.1152 =D16*F16 =G16+H15
17 Year 6 =D16 x 0.0576 =D17*F17 =G17+H16
18
19 Hence accumulated depreciation in year 3 =H14
20